Let x1 > 1 and xn+1 := 2 −

1

xn

for all n ∈ N. Show that {xn} is bounded


and monotone. Find the limit.


1
Expert's answer
2021-11-03T18:25:21-0400

Proof. Since x1 > 1, we have that 1/x1 < 1. This means that x2 =

2 − 1/x1 > 1. We’d like to show that xn > 1 for all n. By induction, we

see that if xk > 1, then xk+1 = 2−1/xk > 1 (easy enough?). At this point

we are bounded above by 2 and below by 1. (Note: We could have made

another arguement that it was bounded below by 0. The bound of 1 is a

little better, but the MCT does not require anything more than bounded.)

Now let’s show that it is monotone (decreasing). To do this, we want to

start by looking at



"0 < (x_1 − 1)^2\\
0 < x^2_1 − 2x_1 + 1\\
2x_1 < x^2_1 + 1\\
2 < x_1 + 1/x_1\\
2 − 1/x_1 < x_1\\
x_2 < x_1"


Ack! That only takes care of the first case. Now we use induction to finish

the monotone.


"x_{k+1} < x_k\\
1/x_{k+1} > 1/x_k\\
−1/x_{k+1} < −1/x_k\\
2 − 1/x_{k+1} < 2 − 1/x_k\\
x_{k+2} < x_{k+1}"


To find the limit, we use the same technique as the examples in the book

and class.


"x_{n+1} = 2 − 1/x_n\\
x = 2 − 1/x\\
x = 1"